# Comparison of marker-less 2D image-based methods for infant pose estimation

### v2

@@ 2.3.2 Statistical tests
When comparing the results with respect to the difference $d_{a}$, we tested the statistical significance of model differences with a paired sample t-test on all individual keypoints in the dataset (or sub-sets of keypoints if indicated). For PCK, where each point is either correct or incorrect, we used a Pearson’s chi-squared test on the frequencies of the outcomes (correct detection or not) to test if the samples for each pose estimation model could come from the same distribution.
@@ 3.1 Variability in human labeling
Figure 2: Difference between two annotators, additionally split by viewing angle.
@@ 3.2 Generic pose estimators
Figure 3: Difference to annotation $d_{a}$ in pixels for different subjects, evaluated on our dataset and grouped by key point
@@
The qualitative result is very close to the difference between human annotators, with the hips having the highest error. But in contrast to the human annotations, the shoulder error is relatively lower, compared to the hip error. Because of the size of the dataset, the differences between the pose estimation models are all statistically significant with $p<0.001$.
The current state of the art model on the COCO dataset, ViTPose-huge, is also the best performing model on our dataset. However, the relative distance to the second best, HRNet-w48, is only 2.4% for PCK@0.1. For high accuracy requirements, e.g., PCK@0.05, the relative distance is only 0.6%, but the difference is still statistically significant with $p=0.0312$. OpenPose, although being an older model, still achieved better results than MediaPipe, which did never achieve a mean error of less than 10 pixels, not even for the clearly defined keypoints like eye and nose.
3.3 Influence of the view angle
Figure 4: Pose estimation errors for the generic pose estimation models, split by viewing angle.
The errors in case of the diagonal view are higher than in the case of the top view. All top/diagonal differences are statistically significant with $p<0.001$. The most pronounced difference is between the hip position estimation errors. In the top view, the difference is not only lower than in the diagonal view, but also lower as compared to the errors of the other keypoints. This is consistent with the results for the difference between human annotators shown in Figure 2.
@@ 3.4 Infant pose estimators
Retraining significantly improved the ViTPose model, with the PCK increasing by 20 percentage points in PCK@0.05 ($p<0.001$). The mean difference to annotation for the hips is decreased by 61%. All other keypoints also see significant improvements (all $p<0.001$).
The other two infant pose estimators Aggpose [35] and AGMA-HRNet48 [36] exhibit different behaviour. Both perform significantly worse than our retrained network ($p<0.001$), with AggPose even performing worse than the non-retrained ViTPose ($p<0.001$). The difference between ViTPose and AGMA-HRNet48 is not statistically significant ($p=0.66$ for $d_{a}$ and $p=0.59$ for PCK@0.05). Regarding the hips, AGMA-HRNet48 improves the accuracy over generic pose estimation, while AggPose is significantly worse (both $p<0.001$).
@@ 3.5 Retraining for individual view angles
There is no significant difference between the model that was trained on both views versus the ones trained on individual views ($p=0.25$ for diagonal, $p=0.54$ for top view), if evaluated against the respective views that were used in training. The results for the respective views that were not used in training, however, are significantly worse ($p<0.001$). This means there is no gain in specializing models to certain viewing angles in our case. However, training on a view that is different from the one used in inference leads to worse performance.
@@
Figure 5: Mean difference from annotation ($d_{a}$) for retrained ViTPose models separately evaluated on diagonal or top view images. Different models have been trained on all, only diagonal view or only top view images, respectively.
@@ 4 Discussion
4.1 Generic pose estimation
The state of the art models ViTPose and HRNet show clear improvements compared to the older model OpenPose. The gain in PCK@0.05 of 5.76 percentage points is small compared to the gain later achieved by retraining, but still statistically significant ($p<0.001$).
In addition to the position accuracy, the reliability of keypoint detection also increases with newer models. A discussion is given in the Supplementary Material 6.
4.2 Influence of the view angle
@@
4.3 Retraining
While the AGMA-HRNet48 does improve pose estimation accuracy of the hip, other keypoints get worse, so there is no statistically significant difference between AGMA-HRNet48 and generic ViTPose ($p=0.66$). It has to be noted that the training data of this model and our training data were both derived from a GMA setting. They differ in the amount of visual clutter in the image (compare with Figure 3 in [36]), but otherwise closely resemble each other. The infant specific retraining improved the HRNet48 by 0.29 and 1.72 percentage points in PCK@0.05 and PCK@0.1 on our dataset, respectively, putting AGMA-HRNet48 on the same performance level as the generic ViTPose. Note, that the AGMA-HRNet48 model is not the best performing one from its original paper [36]. Their DarkPose32 model scored 2.23 and 0.55 percentage points higher than HRNet48 at PCK@0.05 and PCK@0.1, respectively. Since the non-retrained DarkPose32 model performed worse than the non-retrained HRNet48 in [36], we do not expect it to yield a significant improvement over the generic pose estimation on our data either.
To date, there are no publicly available infant datasets suitable for GMA (mostly due to patient data privacy issues; see [30]), so we could not evaluate our model on data from a different setup. This is probably further contributing to the general overfitting situation we identified, because there is no training set available that covers multiple GMA related setups. See Supplementary Material 7 for a comment on a preliminary analysis with the SyRIP dataset [34].
@@ 5 Conclusion
In summary, our study suggests that pose estimation accuracy would benefit from the utility of the top view and the pose estimator being trained on the specific dataset. However, a generic pose estimator should be preferred over a specialized pose estimator trained on a different dataset, if retraining is not feasible. Whether potentially also movement classification accuracy would be improved with more accurate pose estimation remains a question, which needs to be addressed in future work. While the standard GMA method uses a diagonal view for assessment, clinical set-ups aiming to generate data for AI and computer vision approaches should in future consider top-down viewing angles for recording new data.
[…]
6 Keypoint detection rates
@@
Figure 6: Mean difference to annotation $d_{a}$ vs. the ratio of missing detections, based on the keypoint prediction confidence scores. Panel a): overall, panel b): split by viewing angle, solid lines correspond to the top view, dashed lines to the diagonal view.
$m=\frac{1}{N}\sum\limits_{i=1}^{N}H(c-t).$ (2)
The keypoint detection rates for all models are displayed in eFigure 6(a). The result for the human annotation variance (see Section 3.1) was also added for comparison.
Except for MediaPipe, the error decreases when filtering out uncertain points. The increase for MediaPipe comes from the model assigning relatively high scores to the ears, which are among its worst detected points. ViTPose-huge has the lowest error across all missing ratios, MediaPipe the highest. OpenPose achieves a lower error than HRNet for missing ratios over 15%, but this is of limited use, as ViTPose is still better and 15% of points not being detected is insufficient for the applications of automated GMA. Moreover, OpenPose does not even detect all points when thresholding with a score of 0, resulting in the missing ratio never dropping below 2.93%. ViTPose and HRNet, however, produce predictions for every possible point when thresholding at 0. Below the missing ratio for humans (1.41%, because of the ears, see Section 2.1.4), the error of ViTPose saturates, while the error of HRNet increases. This is because ViTPose is assigning low scores to the ear keypoints that are not visible (e.g., due to head turned to the side) and therefore don’t affect the error calculation.
eFigure 6(b) shows the mean difference to annotation in dependence of the missing ratio, like eFigure 6(a), but split by perspective. The performance for the diagonal view is always worse than for the top view. Moreover, while the error for the diagonal view drops faster than for the top view when filtering out uncertain points, they don’t converge to the same level even when filtering out more than 20% of all detected points. We again observe OpenPose not being able to provide estimates for every point, with the diagonal angle having more non-detected keypoints as compared to the top view.
@@
7 SyRIP
The SyRIP dataset [34] was considered, and some preliminary analyses were performed, however, it became obvious that the setting of this dataset is too different from GMA setting, with many more different body positions and much older infants than in our dataset. As in case of the specialized infant pose estimators evaluated on our dataset, our model could not compete with the generic ViTPose on SyRIP. Retraining decreased the PCK@0.05 from 51.53% to 27.65% (73.66% to 46.97% for PCK@0.1), further strengthening our point that specialized infant pose estimators (in this case our own) are overfit and do not generalize well to other datasets.
Acknowledgment

### v3

@@ 2.3.4 Statistical tests
When comparing the results with respect to the difference $d_{a}$, we tested the statistical significance of model differences with a paired sample t-test on all keypoints in the dataset (or sub-sets of keypoints if indicated). For PCK, where each point is either correct or incorrect, we used a Pearson’s chi-squared test on the frequencies of the outcomes (correct detection or not) to test if the samples for each pose estimation model could come from the same distribution.
@@ 3.1 Variability in human labeling
Figure 2: Difference between two annotators, additionally split by viewing angle. Error bars represent confidence intervals of mean (95%).
@@ 3.2 Generic pose estimators
Figure 3: Difference to annotation $d_{a}$ in pixels for different subjects, evaluated on our dataset and grouped by key point. Error bars represent confidence intervals of mean (95%).
@@
The qualitative result is very close to the difference between human annotators, with the hips having the highest error. But in contrast to the human annotations, the shoulder error is relatively lower, compared to the hip error. Because of the size of the dataset, the differences between the pose estimation models are all statistically significant with $p<0.001$ (paired sample t-test on all keypoints).
The current state of the art model on the COCO dataset, ViTPose-huge, is also the best performing model on our dataset. However, the relative distance to the second best, HRNet-w48, is only 2.4% for PCK@0.1. For high accuracy requirements, e.g., PCK@0.05, the relative distance is only 0.6%, but the difference is still statistically significant with $p=0.0312$ (Pearson’s chi-squared test). OpenPose, although being an older model, still achieved better results than MediaPipe, which did never achieve a mean error of less than 10 pixels, not even for the clearly defined keypoints like eye and nose.
@@ 3.3 Infant pose estimators
Retraining significantly improved the ViTPose model, with the PCK increasing by 20 percentage points in PCK@0.05 ($p<0.001$, Pearson’s chi-squared test). The mean difference to annotation for the hips is decreased by 61%. All other keypoints also see significant improvements (all $p<0.001$, individual paired sample t-tests on keypoints).
The other two infant pose estimators Aggpose [41] and AGMA-HRNet48 [42] exhibit different behaviour. Both perform significantly worse than our retrained network ($p<0.001$, paired sample t-test on all keypoints), with AggPose even performing worse than the non-retrained ViTPose ($p<0.001$, paired sample t-test on all keypoints). The difference between ViTPose and AGMA-HRNet48 is not statistically significant ($p=0.66$ for $d_{a}$, paired sample t-test on all keypoints, and $p=0.59$ for PCK@0.05, Pearson’s chi-squared test). Regarding the hips, AGMA-HRNet48 improves the accuracy over generic pose estimation, while AggPose is significantly worse (both $p<0.001$, paired sample t-test on all keypoints).
3.4 Influence of the view angle
Figure 4: Pose estimation errors for the generic pose estimation models, split by viewing angle. Error bars represent confidence intervals of mean (95%).
The errors in case of the diagonal view are higher than in the case of the top view. All top/diagonal differences are statistically significant with $p<0.001$ (paired sample t-test on all keypoints). The most pronounced difference is between the hip position estimation errors. In the top view, the difference is not only lower than in the diagonal view, but also lower as compared to the errors of the other keypoints. This is consistent with the results for the difference between human annotators shown in Figure 2.
@@ 3.5 Retraining for individual view angles
There is no significant difference between the model that was trained on both views versus the ones trained on individual views ($p=0.25$ for diagonal, $p=0.54$ for top view, both paired sample t-tests on all keypoints), if evaluated against the respective views that were used in training. The results for the respective views that were not used in training, however, are significantly worse ($p<0.001$, paired sample t-test on all keypoints). This means there is no gain in specializing models to certain viewing angles in our case. However, training on a view that is different from the one used in inference leads to worse performance.
@@
Figure 5: Mean difference from annotation ($d_{a}$) for retrained ViTPose models separately evaluated on diagonal or top view images. Different models have been trained on all, only diagonal view or only top view images, respectively. Error bars represent confidence intervals of mean (95%).
3.6 Inference speed
Since the focus of our study is not on real-time pose estimation for real time applications, we did not optimize the pose estimators (e.g., architectures and/or meta parameters) for inference speed.
All models can be executed on any recent GPU with at least 4GB of memory (for model complexity please see the respective original publications). Table 2 lists the observed inference speed on a NVIDIA GTX 2080 Ti. Those speeds were obtained by using the demonstration scripts of the respective models and adapting them to process our dataset. The results show that all models have acceptable and reasonable inference speed with the ViTPose being the slowest (but the most accurate), and OpenPose the fastest model.
Table 2: Inference speed in frames per second (fps), on a NVIDIA GTX 2080 Ti, for each included model. This is the average speed over processing the 4500 frames of the dataset.
Model ViTPose-huge HRNet-w48 MediaPipe OpenPose Retrained ViTPose AggPose AGMA-HRNet48
Inference speed 1 fps 1.2 fps 3 fps (CPU) 11.5 fps 1.1 fps 2.7 fps 3 fps
@@ 4 Discussion
Generic pose estimators
The state of the art models ViTPose and HRNet show clear improvements compared to the older model OpenPose. The gain in PCK@0.05 of 5.76 percentage points is small compared to the gain later achieved by retraining, but still statistically significant ($p<0.001$, Pearson’s chi-squared test).
In addition to the position accuracy, the reliability of keypoint detection also increases with newer models (see Figure S1). A discussion is given in the Appendix.
Infant specific pose estimators
While the AGMA-HRNet48 does improve pose estimation accuracy of the hip, other keypoints get worse, so there is no statistically significant difference between AGMA-HRNet48 and generic ViTPose ($p=0.66$, paired sample t-test on all keypoints). It has to be noted that the training data of this model and our training data were both derived from a GMA setting. They differ in the amount of visual clutter in the image (compare with Figure 3 in [42]), but otherwise closely resemble each other. The infant specific retraining improved the HRNet48 by 0.29 and 1.72 percentage points in PCK@0.05 and PCK@0.1 on our dataset, respectively, putting AGMA-HRNet48 on the same performance level as the generic ViTPose. Note, that the AGMA-HRNet48 model is not the best performing one from its original paper [42]. Their DarkPose32 model scored 2.23 and 0.55 percentage points higher than HRNet48 at PCK@0.05 and PCK@0.1, respectively. Since the non-retrained DarkPose32 model performed worse than the non-retrained HRNet48 in [42], we do not expect it to yield a significant improvement over the generic pose estimation on our data either.
To date, there are no publicly available infant datasets suitable for GMA (mostly due to patient data privacy issues; see [35]), so we could not evaluate our model on data from a different setup. This is probably further contributing to the general overfitting situation we identified, because there is no training set available that covers multiple GMA related setups. See the Appendix for a comment on a preliminary analysis with the SyRIP dataset [40].
Influence of viewing angle
@@ 5 Conclusion
We compared four different generic pose estimation frameworks on our dataset of 4500 frames recorded with a multi-view camera setup designed for automated GMA. The current best performing model on COCO test-dev, ViTPose, also achieved the best results on our infant dataset. Retraining the best performing network on our dataset, expectedly, increased the performance, especially on the keypoint that is most difficult to detect, the hips. In comparison with the other infant pose estimation frameworks, we achieved 20 percentage points better PCK@0.05 than the second best model. The other models performed below or on the level of the best generic pose estimator. This suggests that, if possible, one should always retrain on the specific dataset to achieve the best pose estimation accuracy. If retraining is not possible, a state-of-the-art generic pose estimator should be used, unless the dataset is very similar to the training data of a specialized infant pose estimator.
We also compared the pose estimation error between different viewing angles. There is a significant difference in accuracy between the diagonal and top views in both human annotation and neural network pose estimation. Therefore, the results strongly suggest that the top view should be preferred in any setup for automated GMA with only one camera. Since the different viewing angles lead to different accuracy, we tested if specialized pose estimators trained on particular views would improve the accuracy, however, no significant effect was found.
In summary, our study suggests that pose estimation accuracy would benefit from the utility of the top view and the pose estimator being trained on the specific dataset. However, a generic pose estimator should be preferred over a specialized pose estimator trained on a different dataset, if retraining is not feasible. Whether potentially also movement classification accuracy would be improved with more accurate pose estimation remains a question, which needs to be addressed in future work. While the standard GMA method uses a diagonal view for assessment, clinical setups aiming to generate data for AI and computer vision approaches should in future consider top-down viewing angles for recording new data.
Acknowledgements
Author contributions statement
L.J. P.M. and T.K developed the idea and research questions. L.J. performed the analysis, on data annotated by S.F., and wrote the main manuscript text. T.K. supervised the work and edited the manuscript from the first draft onward. D.Z., L.P. and P.M. wrote the sections about the clinical aspects and GMA. S.F., D.Z. S.B. F.W. and P.M. edited the manuscript. All authors read and approved the final manuscript version.
Data availability statement
The authors do not have permission to share the infant video recording data. Derived data, like the annotations and pose estimation data are available from the corresponding author upon reasonable request. The five retrained ViTPose models are published on Zenodo (https://doi.org/10.5281/zenodo.14833182).
Additional information
Competing interests: The authors declare no competing interests.
Appendix A Appendix
A.1 Keypoint detection rates
$m=\frac{1}{N}\sum\limits_{i=1}^{N}H(c-t).$
The keypoint detection rates for all models are displayed in Figure 1(a). The result for the human annotation variance (see Section 3.1) was also added for comparison.
@@ A.1 Keypoint detection rates
Figure S1: Mean difference to annotation $d_{a}$ vs. the ratio of missing detections, based on the keypoint prediction confidence scores. Panel a): overall, panel b): split by viewing angle, solid lines correspond to the top view, dashed lines to the diagonal view.
Except for MediaPipe, the error decreases when filtering out uncertain points. The increase for MediaPipe comes from the model assigning relatively high scores to the ears, which are among its worst detected points. ViTPose-huge has the lowest error across all missing ratios, MediaPipe the highest. OpenPose achieves a lower error than HRNet for missing ratios over 15%, but this is of limited use, as ViTPose is still better and 15% of points not being detected is insufficient for the applications of automated GMA. Moreover, OpenPose does not even detect all points when thresholding with a score of 0, resulting in the missing ratio never dropping below 2.93%. ViTPose and HRNet, however, produce predictions for every possible point when thresholding at 0. Below the missing ratio for humans (1.41%, because of the ears), the error of ViTPose saturates, while the error of HRNet increases. This is because ViTPose is assigning low scores to the ear keypoints that are not visible (e.g., due to head turned to the side) and therefore don’t affect the error calculation.
Figure 1(b) shows the mean difference to annotation in dependence of the missing ratio, like Figure 1(a), but split by perspective. The performance for the diagonal view is always worse than for the top view. Moreover, while the error for the diagonal view drops faster than for the top view when filtering out uncertain points, they don’t converge to the same level even when filtering out more than 20% of all detected points. We again observe OpenPose not being able to provide estimates for every point, with the diagonal angle having more non-detected keypoints as compared to the top view.
@@
A.2 SyRIP
The SyRIP dataset40 was considered, and some preliminary analyses were performed, however, it became obvious that the setting of this dataset is too different from GMA setting, with many more different body positions and much older infants than in our dataset. As in case of the specialized infant pose estimators evaluated on our dataset, our model could not compete with the generic ViTPose on SyRIP. Retraining decreased the PCK@0.05 from 51.53% to 27.65% (73.66% to 46.97% for PCK@0.1), further strengthening our point that specialized infant pose estimators (in this case our own) are overfit and do not generalize well to other datasets.